OpenStudy (anonymous):

If f(x) = 2x and g(x) = x-4, what is the value of f[g(3)]? -6 -2 0 6 2

OpenStudy (anonymous):

i think your suppose to solve for g(x) first by plugging in the 3 to get -1. then you would plug the -1 into f(x)2(-1) to get -2 so your answer would be -2

OpenStudy (anonymous):

candygirl7655 i dont get ur question :(

OpenStudy (campbell_st):

find g(3) so if g(x) = x - 4 then g(3) = 3 -4 g(3) = -1 the find f(-1)

OpenStudy (anonymous):

lol me either its okay :)

OpenStudy (campbell_st):

its just a case of substitution... if g(x) replace x with 3 and solve... then using that answer... in f(x) replace x with -1 and solve

OpenStudy (anonymous):

hey campbell can u help me with mine

OpenStudy (campbell_st):

what's your question

OpenStudy (anonymous):

i have it posted